OpenStudy (anonymous):

The length of a rectangle is 5 more than twice its width, and the area of the rectangle is 88 ft^2 . Find the dimensions of the rectangle

OpenStudy (anonymous):

16 l 5.5 w i think

OpenStudy (anonymous):

\[A=lw\] \[l=2w+5\] \[A=88\] \[88=w(2w+5)\] \[88=2w^2+5w\] \[0=2w^2=5w-88\] \[w=\frac{11}{2}\]

OpenStudy (radar):

Let w = width then from the information given you can state; l=2w+5 and we are given Area as 88 square feet. Area is length times width. w(2w+5)= 88 \[2w ^{2}+5w=88\]\[2w ^{2}+5w-88=0\] That will factor nicely (2w-11) (w+8) w=11/2=5.5 ft. l=16 ft. check for correct area 16 X 5.5= 88 sq ft